OpenStudy (anonymous):

what do you get when you plug in -61776 into cos^-1(x) on your calc? how come i keep getting error domain????

OpenStudy (asnaseer):

the cosine of an angle varies between 1 and -1. so it can never have the value -61776

OpenStudy (zepp):

This ^ :(

OpenStudy (zepp):

I should type a little bit faster..

OpenStudy (asnaseer):

you are asking your calculator to find the angle whose cosine would be -61776 - no such angle exists

OpenStudy (asnaseer):

sorry zepp - didn't mean to step on your toes :)

OpenStudy (anonymous):

ok so im trying to find the largest angle when given sides 22,36, and 44 using cosC= a^2 + b^2 - c^2 / 2 ab

OpenStudy (asnaseer):

kalloohkallay32: you must have made an error in your calculations

OpenStudy (zepp):

The largest side is faced to the hypontenuse, and I didn't getting any error.

OpenStudy (zepp):

didn't get*

OpenStudy (asnaseer):

zepp: the largest angle would indeed be opposite the largest side. but, in this case, we do not have a right-angled triangle, so it is not the hypotenuse.

OpenStudy (anonymous):

common error... use parenthesis when entering in your calculator... cosC= (a^2 + b^2 - c^2 )/ (2 ab) so C = cos^-1 (a^2 + b^2 - c^2 )/ (2 ab)

OpenStudy (asnaseer):

kalloohkallay32: make sure you do the division right in your calculator. you should be calculating:\[\cos(C)=\frac{a^2+b^2-c^2}{2ab}\]

OpenStudy (zepp):

Oh yeah right, thx asnaseer :D

OpenStudy (zepp):

According to my calculator, the largest angle should be 87.17747 degrees.